INAF queue observing, 2013 June 11-12

Observer: F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn
Instrument: LBC -MODS

Summary

At the sunset clear sky. Only some cloud to the horizon.
During the first part of the night some thin cirrus passed through
Seeing around 1".
Strange feature (we assume to be a moth) is visibile in some red frames. (frame 3 of 1.ob, frame 1 of 2.ob, frame 1to4 of 3.ob)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard hz44 Blue grating    
Sommariva_31 1625-2648 Blue grating 2.5h Last exposure has variable seeing
Bedin_29 E NGC6819-bf V-Bessel i-Sloan    

Details

02:32 The TO opens the dome

Start using MODS

03:23 Start with Sommariva standard 003-005
  • acqMODS hz44.acq
  • Send offset dx=-0.615 by hand for standard acquisition

03:34 execMODS hz44.obs 006

03:47 acqMODS n1625_mos.acq

04:05 Redo acquisition using only slit 1 and 4 (due to proper motion of the other stars)
  • Refine the acquisition because star 1 seems not perfectly aligned
  • Looks good now

04:14 execMODS n1625_1.obs 010-011
  • dimm sseing ~1"
  • dimm seeing is bit worse ~1.2"

04:18 execMODS n1625_2.obs 012-013

05:50 Some clouds passed

06:23 execMODS n1625_3.obs 014-015

06:56 Aborted second exposure of Sommariva to switch to LBC

SWITCH LBC

07:12 Run the focus OB

07:28 is not well done (the red image is missing)

07:35 Select another pointing OB from default dir

07:36 Run WT10_332.ob

07:39 Run science OB: N6819_RunIDE_LONG1.ob
  • dofpia
  • 8:00 A strange feature appears on the 1chip of the 3th image in the red channel

08:14 dofpia
  • Run science OB: N6819_RunIDE_LONG2.ob
  • Same strange feature appears on the 2chip of the 1st image (always red channel). Moth in the filter wheel?

08:49 dofpia
  • Run science OB: N6819_RunIDE_LONG3.ob
  • Same strange feature appears in all red images, but the last lbcr.20130612.091337.fits

09:21 We run a script that moves the filter wheels in all positions
  • Stopped at 09:22 for an error
  • Error keeps hang LBC GUI and we are not able to run a new script
  • Reset LBC (lbckill.sh and lbcstart.sh)
  • We created on fly a quick script which moves filters wheel. By this way we try to fix red channel problem, otherwise we have to move telescope to the horizon, close the dome and open the camera and fix the problem, ... (approx 1h, and we are next to the twilight)
  • After script images now look good

09:29 dofpia
  • Run science OB: N6819_RunIDE_LONG4.ob
  • With the observation filter (i-Sloan) the feature is still there

10:03 Slew to the standard star (another attempt to fix the problem with moth)
  • Images seem good, but in this case we have also a shorter exposure time

10:12 Slew back to the target

10:14 dofpia
  • Run again science OB: N6819_RunIDE_LONG3.ob
  • again the black spot, now on the upper part of the LBC field
  • stopped at 10:29 after two exposures for each channel

10:20 Close the dome and try to check the camera

11:16 Open the dome again.

11:23 Back on sky, Try to acquire some flats.
  • Run SkyFlat_Vi_r1921d+12_pa0_5.ob (x3)
  • Run SkyFlat_Vi_r1921d+12_pa180_5.ob (x2)
  • Run SkyFlat_Vi_r1921d+12_pa0_5.ob (x1)
  • Run SkyFlat_Vi_r1921d+12_pa180_5.ob (x0.3)
  • No more features in the red flats.

11:47 Slew to the horizon and close dome.

-- MarcoFumana - 12 Jun 2013
Topic revision: r6 - 15 Jun 2013, MarcoFumana
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback